OpenStudy (anonymous):

solve for x: log6(x+5)+log6(3)=log6(24)

sam (.sam.):

log6(x+5)+log6(3)=log6(24) log6(x+5)+log6(3)-log6(24)=0 using laws of logarithm, \[ log_6\frac{(x+5)(3)}{24}=0\] \[6=\frac{(x+5)(3)}{24}\] can you solve for x now?

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

yeah i'm stuck... do you distribute the 3?

sam (.sam.):

yes

OpenStudy (anonymous):

I end up with 6= 3x+15 all over 24 right?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

x=43?

OpenStudy (anonymous):

thanks

sam (.sam.):

Yes